# Learning with Less: Knowledge Distillation from Large Language Models via Unlabeled Data

### v1

@@ 1 Introduction
To tackle these limitations, smaller models  Liu (2019); Devlin (2018) are often preferred due to their lower resource demands. Nonetheless, smaller models are not as powerful as LLMs Kaplan et al. (2020) and typically require further training for specific tasks using labeled data, as they usually do not have the capacity to capture broad knowledge. Without the guidance of labeled data, self-supervised training on smaller models may lead to suboptimal performance, as these models struggle to generalize across diverse tasks and often fail to learn task-specific features effectively Goyal et al. (2019). This challenge is further hindered by the high cost of obtaining task-related labeled data. While unlabeled data is generally more abundant, it cannot be directly utilized without proper labeling, posing a significant challenge for model training.
One promising approach is to use LLMs to generate pseudo-labels for unlabeled data, which can then be used to train smaller models. This strategy allows smaller models to benefit from the extensive knowledge embedded in the LLM while reducing computational costs. This process can be seen as a form of knowledge distillation Mishra and Sundaram (2021); Zhou et al. (2023). However, this approach presents challenges. Pseudo-labels generated by LLMs may be noisy or unreliable, potentially degrading the performance of the student model. In some cases, the availability of unlabeled data may also be limited, particularly for domain-specific tasks where data collection is challenging Adams et al. (2017); Xiao et al. (2018). Thus, achieving data efficiency is crucial—not only to reduce the impact of noisy pseudo-labels but also to ensure that representative data samples are selected for optimal training.
@@ 4.1 Experimental Settings
Models PubMed-RCT-20k Yahoo! Answers Emotions Arxiv-10 BiosBias Avg. Rank ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 / Teacher 67.09 57.88 65.45 64.12 59.28 48.6 56.02 56.5 75.83 61.32 10 Random 70.17 ± 0.78 61.70 ± 0.72 66.42 ± 0.40 64.95 ± 0.39 65.51 ± 0.23 55.59 ± 0.65 60.86 ± 0.97 60.90 ± 0.77 76.30 ± 0.29 62.49 ± 0.52 5.9 FreeMatch 70.05 ± 0.79 61.37 ± 1.04 66.36 ± 0.56 64.68 ± 0.30 64.90 ± 0.15 54.43 ± 0.43 60.50 ± 0.93 60.41 ± 0.85 76.23 ± 0.09 62.59 ± 0.23 7.8 SoftMatch 70.11 ± 0.19 61.31 ± 0.44 66.56 ± 0.32 64.78 ± 0.52 64.90 ± 0.88 56.01 ± 1.19 61.28 ± 0.41 60.70 ± 0.45 76.26 ± 0.19 62.23 ± 0.44 6.9 CCKD_L 69.67 ± 0.38 61.13 ± 0.54 66.58 ± 0.36 64.89 ± 0.06 63.74 ± 0.48 53.60 ± 0.87 61.15 ± 0.18 60.87 ± 0.25 76.37 ± 0.16 62.66 ± 0.07 6.8 CCKD_T+Reg 70.06 ± 0.43 61.62 ± 0.45 66.76 ± 0.14 64.86 ± 0.12 64.33 ± 0.62 54.15 ± 1.07 61.65 ± 0.85 61.26 ± 1.05 76.74 ± 0.06 65.29 ± 0.22 5.4 UNIXKD 70.87 ± 0.32 62.03 ± 0.61 66.58 ± 0.63 64.89 ± 0.38 65.83 ± 0.18 56.09 ± 0.58 61.73 ± 1.04 61.35 ± 1.07 76.30 ± 0.18 62.58 ± 0.18 4.5 Entropy Score 71.43 ± 0.65 62.53 ± 0.84 66.90 ± 0.18 65.33 ± 0.09 65.08 ± 0.48 55.01 ± 1.13 60.93 ± 0.51 60.21 ± 0.18 76.86 ± 0.39 64.81 ± 0.44 4.6 Lbl2TransformerVec 31.82 23.44 53.74 53.31 42.73 39.67 42.88 41.71 47.78 47.32 11 LLKD 73.46 ± 0.20 65.72 ± 0.22 67.15 ± 0.05 65.57 ± 0.21 67.07 ± 1.07 56.24 ± 1.67 62.85 ± 0.72 62.76 ± 0.60 77.21 ± 0.37 65.44 ± 0.19 1.6 LLKD_w 74.16 ± 0.95 66.44 ± 0.76 67.15 ± 0.30 65.80 ± 0.36 67.44 ± 1.58 56.16 ± 0.95 62.99 ± 0.53 62.94 ± 0.39 76.93 ± 0.07 65.07 ± 0.33 1.5 Improv. 3.82% 6.25% 0.37% 0.72% 2.45% 0.27% 2.04% 2.59% 0.46% 0.23% /
@@
Baselines. We compare our method against four groups of baseline approaches: 1) Thresholding methods, such as FreeMatch (Wang et al., 2023), which uses an adaptive threshold to select samples with high student confidence, and SoftMatch (Chen et al., 2023), which assigns higher weights to samples with greater student confidence. 2) Knowledge distillation methods, which focus on filtering noisy pseudo-labels based on the teacher or selecting informative samples based on the student. For the first type, we evaluate CCKD_L Mishra and Sundaram (2021), which weights samples according to the teacher’s probability, and Entropy Score (Lang et al., 2022), which selects samples with the lowest teacher entropy, indicating high teacher confidence. For the second type, we include CCKD_T+Reg Mishra and Sundaram (2021), which uses a threshold to select challenging samples for the student, and UNIXKD(Xu et al., 2023), which selects samples with the highest student uncertainty. 3) Unsupervised text classification method: Lbl2TransformerVec(Schopf et al., 2022), which predicts labels based on the similarity between the embeddings of input text and label words. 4) Basic baselines: Random, where a subset of samples is randomly selected from each batch, and Teacher, where predictions are made directly using the teacher model.
Implementation details. For the teacher model, we use LLaMA Touvron et al. (2023), an open-source LLM that has demonstrated strong performance in various applications. For the student model, we adopt RoBERTa Liu (2019). To ensure a fair comparison, all baseline models use the same pseudo-labels as our model, and the baseline models use RoBERTa as the backbone model except for the Lbl2TransformerVec. Model performance on the text classification task is evaluated using Accuracy (ACC) and Macro-F1 scores, the data efficiency is evaluated based on the total used training number. We provide a parameter analysis in Section 5. Additional details are given in Section A.4 in the Appendix.
@@
Models PubMed-RCT-20k Yahoo! Answers Emotions Arxiv-10 BiosBias FreeMatch 469,785 (87%) 266,314 (88.8%) 240,388 (72.1%) 413,283 (86.1%) 356,578 (92.4%) CCKD_T+Reg 214,967 (39.8%) 163,259 (54.4%) 159,289 (47.8%) 418,463 (87.2%) 147,427 (38.2%) UNIXKD 50,634 (9.4%) 150,000 (50%) 93,780 (28.1%) 45,000 (9.4%) 265,452 (68.8%) Entropy Score 270,045 (50%) 209,937 (70%) 166,715 (50%) 335,874 (70%) 270,262 (70%) LLKD 19,828 (3.7%) 64,697 (21.6%) 80,725 (24.2%) 188,540 (39.3%) 69,993 (18.1%) w/o SU 366,423 (67.8%) 232,643 (77.6%) 197,757 (59.3%) 362,421 (75.5%) 307,795 (79.7%) w/o TC 71,662 (13.3%) 116,661 (38.9%) 150,091 (45%) 288,185 (60.1%) 134,509 (34.8%)
@@ 4.2 Classification Performance Comparison
Our key observations are as follows: 1) As shown in Table 1, our model consistently outperforms all baseline methods, achieving a significant relative improvement of 6.25% on the Pubmed-rct-20k dataset in terms of F1 score. The weighted version generally performs better, indicating that leveraging teacher confidence and student uncertainty to prioritize selected samples further enhances overall model performance. 2) Generally, we observe that directly using the teacher model performs worse than our method, as well as all other baselines (except for Lbl2TransformerVec), which fine-tune the student model using pseudo-labels. This demonstrates that the student model not only effectively learns from the teacher but also achieves superior results. These findings suggest that, with proper tuning, the student model can deliver better performance while maintaining much lower computational costs. 3) Additionally, the similarity-based method, Lbl2TransformerVec, exhibit the weakest performance, highlighting that relying solely on text-label similarity is insufficient for effective classification.
@@ 4.3 Ablation Study
In this subsection, we conduct an ablation study to evaluate the effectiveness of each component in our method, including the pseudo-label, teacher confidence, and student uncertainty. The results are shown in Figure 3. We use “w/o TC” to denote the model without using the teacher confidence to select samples, relying solely on the student uncertainty threshold. Similarly, “w/o SU” denotes the model without using the student uncertainty threshold to select, while “w/o TC+SU” represents the model without any data selection. The figure clearly shows that our model consistently outperforms the ablated versions across all datasets, highlighting the importance of each component. Notably, when no data selection is performed, the model exhibits the worst performance, further validating the critical role of our data selection strategy in enhancing the model performance.
@@ 4.4 Data Efficiency
In this subsection, we evaluate data efficiency by presenting the total number of samples seen and the percentage of original total samples seen in each run. For example, the original training size of Arxiv-10 is 79,790 and we set the epoch to be 6, so the student model sees a total of 79,790 x 6 = 478,740 samples without any data selection. All methods have the same original total seen samples. The results of various data selection method are shown in Table 2. We also include results from our ablated versions to provide deeper insights into our approach. For UNIXKD and Entropy Score which have fixed selection ratio, we experimented with ratios of {10%, 30%, 50%, 70%, 90%} and used the ratio that achieved the best validation performance. Since SoftMatch and CCKD_L weight samples instead of selecting, they use the full set of original samples. Generally, the results indicate that our method consistently outperforms others in both effectiveness and data efficiency. In most cases , our approach requires selecting only about 20% of the training samples. Notably, on the PubMed-RCT-20k dataset, we use as little as 3.7% of the training samples while achieving a significant relative improvement of 6.25%, as shown in Table 1. Although the UNIXKD is trained with less data on the Arxiv-10 dataset, our method has large performance improvement, indicating UNIXKD does not select sufficient informative data.
@@ 4.5 Choice of Teacher Model
Models ACC F1 Teacher 59.5 59.5 Random 61.56 ± 0.85 61.12 ± 0.59 FreeMatch 61.23 ± 0.54 60.64 ± 1.04 SoftMatch 60.94 ± 0.32 60.57 ± 0.76 CCKD_L 61.18 ± 0.49 60.54 ± 0.53 CCKD_T+Reg 61.22 ± 0.30 60.68 ± 0.32 UNIXKD 61.40 ± 0.60 61.08 ± 0.83 Entropy Score 61.23 ± 0.66 60.59 ± 0.53 Lbl2TransformerVec 42.88 41.71 LLKD 62.84 ± 0.68 62.67 ± 0.17
@@ A.1 Prompts
We provide more details about the prompts used for the teacher and student models. The teacher prompt template is shown in Table 4 and consists of several components: a system prompt with task instructions, examples from the validation set demonstrating the task, the input text, and the output label. The system prompt for each dataset is listed in Table 7. For the student model, we use the template “<Input text>. It is [MASK]".
@@ A.3 Datasets
  • PubMed-RCT-20k [^2] https://github.com/Franck-Dernoncourt/pubmed-rct?tab=readme-ov-file(Dernoncourt and Lee, 2017) is a sequential sentence classification dataset derived from PubMed abstracts, where each sentence is labeled with roles such as background, objective, method, result, or conclusion. The dataset contains 20k abstracts and has fixed train, validation, and test splits.
  • Yahoo! Answers [^3] https://github.com/LC-John/Yahoo-Answers-Topic-Classification-Dataset/tree/master (Zhang et al., 2015) is a topic classification dataset collected from the Yahoo Answers platform consisting of 10 classes. Due to computational limitations, we randomly select a subset: 10k samples from each class for training, and 6k samples from each class for validation.
@@
  • BiosBias [^6] https://huggingface.co/datasets/LabHC/bias_in_bios?row=46 (De-Arteaga et al., 2019) contains textual biographies aimed at predicting professional occupations with 28 classes. This dataset includes predefined splits for training, validation, and testing.
@@
Datasets #Train #Val #Test #Class Labels PubMed-RCT-20k 180,030 30,212 30,135 5 BACKGROUND, OBJECTIVE, METHODS, RESULTS, CONCLUSIONS Yahoo! Answers 99,970 60,000 60,000 10 society, science, health, education, computer, sports, business, entertainment, relationship, politics Emotions 333,431 41,680 41,680 6 sadness, joy, love, anger, fear, surprise Arxiv-10 79,970 10,000 10,000 10 astro-ph, cond-mat, cs, eess, hep-ph, hep-th, math, physics, quant-ph, stat BiosBias 257,394 39,600 99,100 28 accountant,architect, attorney, doctor, comedian, composer, dentist, nutritionist, dj, filmmaker, designer,journalist, model, nurse, painter, assistant, pastor, trainer, photographer, physician, poet, professor, psychologist, rapper, programmer, surgeon,teacher, instructor

### v2

@@ 1 Introduction
To tackle these limitations, smaller models  Liu (2019); Devlin (2018); Wang et al. (2024) are often preferred due to their lower resource demands. Nonetheless, smaller models are not as powerful as LLMs Kaplan et al. (2020) and typically require further training for specific tasks using labeled data, as they usually do not have the capacity to capture broad knowledge. Without the guidance of labeled data, self-supervised training on smaller models may lead to suboptimal performance, as these models struggle to generalize across diverse tasks and often fail to learn task-specific features effectively Goyal et al. (2019). This challenge is further hindered by the high cost of obtaining task-related labeled data. While unlabeled data is generally more abundant, it cannot be directly utilized without proper labeling, posing a significant challenge for model training.
One promising approach is to use LLMs to generate pseudo-labels for unlabeled data, which can then be used to train smaller models. This strategy allows smaller models to benefit from the extensive knowledge embedded in the LLM while reducing computational costs. This process can be seen as a form of knowledge distillation Mishra and Sundaram (2021); Zhou et al. (2023); Kontonis et al. (2024); Iliopoulos et al. (2022). However, this approach presents challenges. Pseudo-labels generated by LLMs may be noisy or unreliable, potentially degrading the performance of the student model. Thus, achieving data efficiency is crucial—not only to reduce the impact of noisy pseudo-labels but also to ensure that representative data samples are selected for optimal training.
@@
Our contributions are summarized as follows: 1) We introduce a knowledge distillation approach that leverages unlabeled data while requiring fewer computational resources. 2) We propose a dynamic data selection method LLKD that identifies high-quality samples, improving data efficiency. 3) Extensive experiments demonstrate that LLKD achieves superior text classification performance and enhanced data efficiency.
@@ 4.1 Experimental Settings
Models PubMed-RCT-20k Yahoo! Answers Emotions Arxiv-10 BiosBias Avg. Rank ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 / Teacher-ZS 53.67 44.77 57.84 58.91 55.15 45.64 53.45 53.08 68.76 51.9 11.9 Teacher 67.09 57.88 65.45 64.12 59.28 48.6 56.02 56.5 75.83 61.32 10.4 Random 70.35 ± 0.73 60.70 ± 1.23 65.91 ± 0.31 64.58 ± 0.42 64.60 ± 1.40 54.35 ± 0.68 60.35 ± 1.56 60.19 ± 1.52 76.11 ± 0.47 61.18 ± 1.35 7.1 No_DS 68.99 ± 0.88 60.40 ± 0.47 65.38 ± 0.92 64.06 ± 0.75 64.55 ± 1.53 54.79 ± 2.34 60.57 ± 0.54 60.16 ± 0.53 75.96 ± 0.46 62.09 ± 0.80 8.2 FreeMatch 69.17 ± 0.39 60.31 ± 1.17 65.98 ± 0.41 64.26 ± 0.12 64.81 ± 0.34 52.68 ± 0.94 60.58 ± 0.91 60.36 ± 1.17 75.63 ± 0.32 61.37 ± 0.76 8.1 SoftMatch 69.63 ± 0.97 60.56 ± 1.51 66.47 ± 0.16 64.70 ± 0.84 65.48 ± 0.46 54.64 ± 1.03 61.01 ± 0.81 59.62 ± 0.78 76.09 ± 0.35 61.20 ± 1.20 6.4 CCKD_L 69.43 ± 0.51 60.61 ± 1.11 65.93 ± 0.32 64.27 ± 0.41 63.19 ± 1.31 52.70 ± 1.35 61.15 ± 0.18 60.10 ± 1.25 76.41 ± 0.17 61.84 ± 0.80 7.3 CCKD_T+Reg 69.89 ± 0.69 60.97 ± 0.87 66.36 ± 0.51 64.53 ± 0.05 64.06 ± 1.32 52.89 ± 0.59 60.75 ± 1.62 60.74 ± 1.51 76.43 ± 0.12 64.04 ± 1.01 5.8 UNIXKD 70.47 ± 0.97 61.65 ± 0.59 66.56 ± 0.70 64.49 ± 1.05 65.29 ± 0.45 55.64 ± 1.15 61.37 ± 1.31 61.35 ± 1.07 75.94 ± 0.51 61.86 ± 0.85 4.6 Entropy Score 71.16 ± 0.91 62.53 ± 0.84 66.61 ± 0.29 65.08 ± 0.53 64.62 ± 0.78 55.42 ± 0.83 60.86 ± 0.57 59.84 ± 0.45 76.62 ± 0.19 64.18 ± 0.48 4.2 Lbl2TransformerVec 31.82 23.44 53.74 53.31 42.73 39.67 42.88 41.71 47.78 47.32 12.9 LLKD 73.46 ± 0.20 65.72 ± 0.22 66.82 ± 0.49 65.5 ± 0.06 67.04 ± 1.12 55.42 ± 2.62 62.90 ± 0.99 62.18 ± 1.08 76.94 ± 0.13 64.63 ± 0.30 1.7 LLKD_w 74.07 ± 0.94 66.17 ± 1.22 66.83 ± 0.12 65.48 ± 0.46 67.03 ± 1.65 56.24 ± 1.04 62.99 ± 0.33 62.90 ± 0.15 76.77 ± 0.16 64.46 ± 0.43 1.4 Improv. 4.09% 5.82% 0.33% 0.80% 2.38% 1.07% 2.64% 2.53% 0.42% 0.70% /
@@
Baselines. We compare our method against four groups of baseline approaches: 1) Thresholding methods, such as FreeMatch (Wang et al., 2023), which uses an adaptive threshold to select samples with high student confidence, and SoftMatch (Chen et al., 2023), which assigns higher weights to samples with greater student confidence. 2) Knowledge distillation methods, which focus on filtering noisy pseudo-labels based on the teacher or selecting informative samples based on the student. For the first type, we evaluate CCKD_L Mishra and Sundaram (2021), which weights samples according to the teacher’s probability, and Entropy Score (Lang et al., 2022), which selects samples with the lowest teacher entropy, indicating high teacher confidence. For the second type, we include CCKD_T+Reg Mishra and Sundaram (2021), which uses a threshold to select challenging samples for the student, and UNIXKD(Xu et al., 2023), which selects samples with the highest student uncertainty. 3) Unsupervised text classification method: Lbl2TransformerVec(Schopf et al., 2022), which predicts labels based on the similarity between the embeddings of input text and label words. 4) Basic baselines: Random, which selects a random subset of samples from each batch; Teacher, which generates predictions directly using the teacher model with few-shot examples; Teacher-ZS, which generates predictions using the teacher model without examples (zero-shot); and No_DS, where the student model is trained without data selection.
Implementation details. For the teacher model, we use LLaMA Touvron et al. (2023), an open-source LLM that has demonstrated strong performance in various applications. For the student model, we adopt RoBERTa Liu (2019). To ensure a fair comparison, all baseline models use the same pseudo-labels as our model, and the baseline models use RoBERTa as the backbone model except for the Lbl2TransformerVec. Model performance on the text classification task is evaluated using Accuracy (ACC) and Macro-F1 scores, the data efficiency is evaluated based on the total used training number. We provide a parameter analysis in Section 5. Additional details are given in Section A.1 and Section A.4 in the Appendix.
@@
Models PubMed-RCT-20k Yahoo! Answers Emotions Arxiv-10 BiosBias FreeMatch 469,218 (87%) 266,314 (88.8%) 240,144 (72%) 415,222 (86.5%) 356,578 (92.4%) CCKD_T+Reg 214,967 (39.8%) 118,502 (39.5%) 242,133(72.6%) 449,526(93.9%) 195,307 (50.6%) UNIXKD 151,875 (28.1%) 262,500 (87.5%) 93,780 (28.1%) 45,000 (9.4%) 337,876 (87.5%) Entropy Score 270,045 (50%) 209,937 (70%) 233,401 (70%) 335,874 (70%) 347,481 (90%) LLKD 19,828 (3.7%) 101,396 (33.8%) 80,725 (24.2%) 161,008 (33.6%) 39,423 (10.2%) w/o SU 366,423 (67.8%) 238,107 (79.4%) 197,757 (59.3%) 376,080 (78.4%) 315,512 (81.7%) w/o TC 71,662 (13.3%) 155,547 (51.9%) 150,091 (45%) 242,105 (50.5%) 83,259 (21.6%)
@@ 4.2 Classification Performance Comparison
Our key observations are as follows: 1) As shown in Table 1, our model consistently outperforms all baseline methods, achieving a significant relative improvement of 5.82% on the Pubmed-rct-20k dataset in terms of F1 score. The weighted version generally performs better, indicating that leveraging teacher confidence and student uncertainty to prioritize selected samples further enhances overall model performance. 2) Generally, we observe that directly using the teacher model performs worse than our method, as well as all other baselines (except for Lbl2TransformerVec), which fine-tune the student model using pseudo-labels. This demonstrates that the student model not only effectively learns from the teacher but also achieves superior results. These findings suggest that, with proper tuning, the student model can deliver better performance while maintaining much lower computational costs. 3) The teacher model (Teacher) with few-shot examples generally outperforms Teacher-ZS, demonstrating the effectiveness of incorporating few-shot examples. 4) Additionally, the similarity-based method, Lbl2TransformerVec, exhibit the weakest performance, highlighting that relying solely on text-label similarity is insufficient for effective classification.
@@ 4.3 Ablation Study
In this subsection, we conduct an ablation study to evaluate the effectiveness of each component in our method, including the data selection, teacher confidence, and student uncertainty. The results are shown in Figure 3. We use “w/o TC” to denote the model without using the teacher confidence to select samples, relying solely on the student uncertainty threshold. Similarly, “w/o SU” denotes the model without using the student uncertainty threshold to select, while “w/o TC+SU” represents the model without any data selection. Note that “w/o TC+SU” is identical to the No_DS in Table 1. The figure clearly shows that our model consistently outperforms the ablated versions across all datasets, highlighting the importance of each component. Notably, when no data selection is performed, the model exhibits the worst performance, further validating the critical role of our data selection strategy in enhancing the model performance.
Additionally, to assess the necessity of utilizing unlabeled data when labeled data is limited, we conduct an analysis, with further details provided in Section A.5 in the Appendix. The results in Table 6 indicate that training with a limited number of true labels yields significantly worse performance compared to the LLKD model trained with unlabeled data. This finding highlights the importance of leveraging unlabeled data when labeled data is limited.
@@ 4.4 Data Efficiency
In this subsection, we evaluate data efficiency by presenting the total number of samples seen and the percentage of original total samples seen in each run. For example, the original training size of Arxiv-10 is 79,790 and we set the epoch to be 6, so the student model sees a total of 79,970 x 6 = 479,820 samples without any data selection. All methods have the same original total seen samples. The results of various data selection method are shown in Table 2. We also include results from our ablated versions to provide deeper insights into our approach. For UNIXKD and Entropy Score which have fixed selection ratio, we experimented with ratios of {10%, 30%, 50%, 70%, 90%} and used the ratio that achieved the best validation performance. Since SoftMatch and CCKD_L weight samples instead of selecting, they use the full set of original samples. Generally, the results indicate that our method consistently outperforms others in both effectiveness and data efficiency. In most cases, our approach requires selecting less than 25% of the training samples. Notably, on the PubMed-RCT-20k dataset, we use as little as 3.7% of the training samples while achieving a significant relative improvement of 5.82%, as shown in Table 1. Although the UNIXKD is trained with less data on the Arxiv-10 dataset, our method has large performance improvement, indicating UNIXKD does not select sufficient informative data.
@@ 4.5 Choice of Teacher Model
Models ACC F1 Teacher 59.5 59.5 Random 61.56 ± 0.85 60.86 ± 1.03 No_DS 60.87 ± 0.66 59.72 ± 0.67 FreeMatch 60.98 ± 0.79 60.30 ± 1.35 SoftMatch 61.55 ± 1.06 60.96 ± 0.91 CCKD_L 60.68 ± 0.46 59.98 ± 0.65 CCKD_T+Reg 61.17 ± 0.64 59.44 ± 1.82 UNIXKD 61.40 ± 0.60 61.08 ± 0.83 Entropy Score 60.34 ± 0.95 59.53 ± 1.17 Lbl2TransformerVec 42.88 41.71 LLKD 62.84 ± 0.68 62.09 ± 0.71
@@ 6 Limitations
7 Acknowledgements
This research is supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, the Army Research Office (ARO) under grant number W911NF-21-1-0198, Amazon Faculty Award, JP Morgan Faculty Award, Meta, Microsoft and SNAP.
@@ A.1 Prompts
We provide more details about the prompts used for the teacher and student models. The teacher prompt template is shown in Table 4 and consists of several components: a system prompt with task instructions, examples from the validation set demonstrating the task, the input text, and the output label. The system prompt for each dataset is listed in Table 8. To determine the number of few-shot examples, we tried number of 3, 5, 10 and chose the one with the best validation result. Specifically, we set the few-shot number as follows: PubMed-RCT-20k (5), Yahoo! Answers (5), Emotions (5), Arxiv-10 (3), and BiosBias (10). For the student model, we use the template “<Input text>. It is [MASK]".
@@ A.3 Datasets
  • PubMed-RCT-20k [^2] https://github.com/Franck-Dernoncourt/pubmed-rct?tab=readme-ov-file(Dernoncourt and Lee, 2017) is a sequential sentence classification dataset derived from PubMed abstracts, where each sentence is labeled with roles such as background, objective, method, result, or conclusion. The dataset contains 20k abstracts and has fixed train, validation and test splits.
  • Yahoo! Answers [^3] https://github.com/LC-John/Yahoo-Answers-Topic-Classification-Dataset/tree/master (Zhang et al., 2015) is a topic classification dataset collected from the Yahoo Answers platform consisting of 10 classes. Due to computational limitations, we randomly select a subset: 10k samples from each class for training, and 6k samples from each class for validation and test.
@@
  • BiosBias [^6] https://huggingface.co/datasets/LabHC/bias_in_bios?row=46 (De-Arteaga et al., 2019) contains textual biographies aimed at predicting professional occupations with 28 classes. This dataset includes predefined splits for training, validation and testing.
By following current works (Kontonis et al., 2024; Iliopoulos et al., 2022), we randomly select 500 labeled samples as the validation set from the original validation set.
Table 6: ACC and Macro-F1 (%) of the text classification. Average results with standard deviation based on 3 seeds are reported.
Models Emotions Arxiv-10 ACC F1 ACC F1 True Label 47.93 ± 2.32 23.10 ± 5.24 53.33 ± 2.72 53.82 ± 1.80 LLKD 67.04 ± 1.12 55.42 ± 2.62 62.90 ± 0.99 62.18 ± 1.08
@@
Datasets #Train #Val #Test #Class Labels PubMed-RCT-20k 180,030 500 30,135 5 BACKGROUND, OBJECTIVE, METHODS, RESULTS, CONCLUSIONS Yahoo! Answers 99,970 500 60,000 10 society, science, health, education, computer, sports, business, entertainment, relationship, politics Emotions 333,431 500 41,680 6 sadness, joy, love, anger, fear, surprise Arxiv-10 79,970 500 10,000 10 astro-ph, cond-mat, cs, eess, hep-ph, hep-th, math, physics, quant-ph, stat BiosBias 257,394 500 99,100 28 accountant,architect, attorney, doctor, comedian, composer, dentist, nutritionist, dj, filmmaker, designer,journalist, model, nurse, painter, assistant, pastor, trainer, photographer, physician, poet, professor, psychologist, rapper, programmer, surgeon,teacher, instructor